# Signatures of Rapidly Rotating Stars with Chemically Homogeneous Evolution in the First Galaxies

### v1

@@ 4 Summary and Discussion
We explore the contribution of rapidly rotating stars under chemically homogeneous evolution (CHE) to the rest-frame UV ($\lambda\sim 1500\ \rm\AA$) emission of the first galaxies. By incorporating CHE into a simple population synthesis model under the assumptions of black-body spectrum and negligible nebular emission, we predict the UV emission efficiency $\eta_{\rm UV}\propto L_{\nu}(\rm UV)/SFR$ as a function of the mass fraction $f_{\rm CHE}$ and minimum mass $m_{\star,\min}^{\rm CHE}$ of CHE stars with two power-law initial mass function (IMF) models ($dN/dm_{\star}\propto m_{\star}^{-\alpha}$, $\alpha=2.35$ and 1). We then adopt the model in Shen et al. (2023) to calculate the relevant UV luminosity function (UVLF) and make comparisons with observations.
@@
We find that CHE can significantly enhance the rest-frame UV emission of star-forming galaxies, as efficient as alternative scenarios of enhanced star formation efficiency and top-heavy IMFs. Compared with the canonical value from local calibrations (Madau & Dickinson, 2014), $\eta_{\rm UV}$ is increased by a factor of $\sim 3-6$ for $f_{\rm CHE}\gtrsim 0.5$ and $m_{\star,\min}^{\rm CHE}\sim 2-10\ \rm M_{\odot}$ with a Salpeter IMF ($\alpha=2.35$, $m_{\star}<300\ \rm M_{\odot}$) at the typical star-formation timescale $t_{\rm SF}\sim 100$ Myr. The effect of CHE is generally stronger for larger $t_{\rm SF}$ or IMF slope $\alpha$. The enhanced UV emission from CHE can reproduce the UVLFs inferred from observations at $z\sim 12-16$ with less extreme values of star formation efficiency and stochastic variability of UV luminosity, within the range of current theoretical predictions.
@@
In addition to boosting the UV emission (as well as ionizing power, see e.g., Sibony et al., 2022) of galaxies, populations of chemically homogeneous stars may have consequences for the nature of transients and for stellar nucleosynthesis. If homogeneous evolution is triggered by fast rotation, those stars (at least the massive ones) may produce long soft gamma ray bursts (GRBs), via the collapsar mechanism (e.g., Yoon & Langer, 2005; Cantiello et al., 2007; Yoon et al., 2012). With the extension to high redshifts considered here, GRBs may then also originate in the first galaxies (e.g., Bromm & Loeb, 2006), and their associated bright afterglows could act as background sources to probe the early intergalactic medium with absorption spectroscopy (Wang et al., 2012). The latter would be an ideal target for the upcoming suite of extremely large telescopes (ELTs). CHE may also play important roles in binary stellar evolution, regulating the properties of binary products, such as Wolf-Rayet stars, X-ray binaries, and binary black hole mergers (e.g., Mandel & de Mink, 2016; de Mink & Mandel, 2016; Marchant et al., 2017; du Buisson et al., 2020; Riley et al., 2021, Dall’Amico et al. in prep.). Moreover, those stars, if they lose mass, may enrich their surroundings in helium, hydrogen-burning products, and possibly also some helium-burning products, while being depleted in lithium. This material could furthermore exhibit strong carbon over-abundances, thus leading to the formation of next-generation carbon-enhanced metal-poor stars (Liu et al., 2021; Jeena et al., 2023). All these aspects would deserve to be studied in future works, illustrating the various manifestations of rapid rotation in the early universe.

### v2

@@ 4 Summary and Discussion
We explore the contribution of rapidly rotating stars under chemically homogeneous evolution (CHE) to the rest-frame UV ($\lambda\sim 1500~{}\rm\AA$) emission of the first galaxies. By incorporating CHE into a simple population synthesis model under the assumptions of black-body spectrum and negligible nebular emission, we predict the UV emission efficiency $\eta_{\rm UV}\propto L_{\nu}(\rm UV)/SFR$ as a function of the mass fraction $f_{\rm CHE}$ and minimum mass $m_{\star,\min}^{\rm CHE}$ of CHE stars with two power-law initial mass function (IMF) models ($dN/dm_{\star}\propto m_{\star}^{-\alpha}$, $\alpha=2.35$ and 1, $m_{\star}<300~{}\rm M_{\odot}$). We then adopt the model in Shen et al. (2023) to calculate the relevant UV luminosity function (UVLF).
@@
We find that CHE can significantly enhance the rest-frame UV emission of star-forming galaxies, as efficient as alternative scenarios of enhanced star formation efficiency and top-heavy IMFs. Compared with the canonical value from local calibrations (Madau & Dickinson, 2014), $\eta_{\rm UV}$ is increased by a factor of $\sim 3-6$ for $f_{\rm CHE}\gtrsim 0.5$ and $m_{\star,\min}^{\rm CHE}\sim 2-10~{}\rm M_{\odot}$ with a Salpeter IMF ($\alpha=2.35$) at the typical star-formation timescale $t_{\rm SF}\sim 100$ Myr. The effect of CHE is generally stronger for larger $t_{\rm SF}$ or IMF slope $\alpha$. The enhanced UV emission from CHE can reproduce the UVLFs inferred from observations at $z\sim 12-16$ with less extreme values of star formation efficiency (SFE) and stochastic variability of UV luminosity, within the range of current theoretical predictions. Taking into account nebular emission can further enhance the effect of CHE by up to a factor of $\sim 2$ for $Z\gtrsim 0.0004$ (see Appendix B).
Moreover, compared with top-heavy IMFs, a potential advantage of the CHE scenario is that the enhancement of SN feedback is likely weaker, making it less vulnerable to the global suppression of star formation by such feedback (Jeong et al., 2024). We estimate the (IMF-averaged) SN energy per unit stellar mass $\mathcal{E}_{\rm SN}$ for our pure CHE and non-rotating populations using the SN models in Goswami et al. (2022), assuming that the SN explosion of a star undergoing CHE is identical to that of the (pre-SN) helium core with the same mass (ignoring mass loss) [^5] We have ignored two competing effects possibly associated with CHE: the enhancement of SN energy by fast rotation (e.g., Burrows et al., 2007; Maeder & Meynet, 2012; Chen, 2015; Fujibayashi et al., 2023), and the reduction of core mass by strong Wolf-Rayet-like winds during post MS or near the end of MS (e.g., Yoon et al., 2012; Kubátová et al., 2019; Riley et al., 2021; Umeda & Nagele, 2024). Taking them into account in a detailed investigation of the SN feedback from CHE stars is deferred to future work.. In the metal-poor regime ($Z\lesssim 0.01$) of high-$z$ galaxies, $\mathcal{E}_{\rm SN}$ is enhanced by a factor of $\sim 25-34$ in the top-heavy ($\alpha=1$) model of non-rotating stars compared with the canonical case of Salpeter IMF ($\alpha=2.35$). On the other hand, in the optimal CHE model with $\alpha=2.35$ and $m_{\star,\min}^{\rm CHE}=4~{}\rm M_{\odot}$, $\mathcal{E}_{\rm SN}$ is only boosted by a factor of $\sim 8-11$, comparable to the increase of $\eta_{\rm UV}$. This means that the effects of CHE on UV emission and SN feedback can be similar to those of increasing the SFE. However, unlike the latter, CHE is not limited by the baryon mass budget.
Besides, stellar UV spectra become harder under CHE (see also Sibony et al., 2022, Liu et al. in prep.) which may play a role in the decrease of UV continuum slope $\beta$ (from the fit $L_{\lambda}\propto\lambda^{\beta}$ for $\lambda\sim 1270-2600\ \rm\AA$) with redshift and extremely blue galaxies with $\beta\lesssim-2.6$ in observations (e.g., Dottorini et al., 2024). In our top-heavy (Salpeter) IMF models, for CHE stars with $m_{\star,\min}^{\rm CHE}=1.7\ \rm M_{\odot}$, we obtain $\beta\sim-3$ ($-2.4$) and $-2.3$ ($-2$) at 10 and 100 Myr after the starburst, respectively, while non-rotating stars give larger values, $\beta\sim-2.7$ ($-2.2$) and $-1.2$ ($-1$), under the same condition [^6] Note that these results do not consider nebular enmission and treat stars as black bodies. According to the tlusty stellar atmosphere models, the spectral shape of UV continuum at $\lambda\sim 1270-2600\ \rm\AA$ does not deviate much from that of black bodies for hot stars. Besides, the nebular contributions can be small in extremely blue galaxies due to high escape fractions and/or bursty star formation (Dottorini et al., 2024)..
@@
Although this work focuses on the UV continuum, stellar and nebular lines are also affected by CHE and fast rotation in general (e.g., Kubátová et al., 2019; Newman et al., 2025). In addition to boosting the UV emission (as well as ionizing power, see, e.g., Sibony et al., 2022, Liu et al. in prep.) of galaxies, populations of chemically homogeneous stars may have consequences for the nature of transients and for stellar nucleosynthesis. If homogeneous evolution is triggered by fast rotation, those stars (at least the massive ones) may produce long soft gamma ray bursts (GRBs), via the collapsar mechanism (e.g., Yoon & Langer, 2005; Yoon et al., 2006; Cantiello et al., 2007; Yoon et al., 2012). With the extension to high redshifts considered here, GRBs may then also originate in the first galaxies (e.g., Bromm & Loeb, 2006), and their associated bright afterglows could act as background sources to probe the early intergalactic medium with absorption spectroscopy (Wang et al., 2012). The latter would be an ideal target for the upcoming suite of extremely large telescopes (ELTs). CHE can play important roles in post-MS and binary stellar evolution, regulating the properties of evolved stars and binary products, such as Wolf-Rayet stars, X-ray binaries, and binary black hole mergers (e.g., Eldridge & Stanway, 2012; Yoon et al., 2012; Mandel & de Mink, 2016; de Mink & Mandel, 2016; Marchant et al., 2017; Kubátová et al., 2019; du Buisson et al., 2020; Riley et al., 2021; Umeda & Nagele, 2024; Dall’Amico et al., 2025). Moreover, those stars, if they lose mass, may enrich their surroundings in helium, hydrogen-burning products, and possibly also some helium-burning products, while being depleted in lithium. This material could furthermore exhibit strong carbon over-abundances, thus leading to the formation of next-generation carbon-enhanced metal-poor stars (Liu et al., 2021; Jeena et al., 2023). All these aspects would deserve to be explored in future studies, illustrating the various manifestations of rapid rotation in the early universe.
@@ Appendix A SFR-UV conversion factors for different star-formation timescales
Figs. 5 and 6 show the results of $\eta_{\rm UV}$ for $t_{\rm SF}=10$ and 1000 Myr, respectively, complementing the results for $t_{\rm SF}=100$ Myr in Fig. 2.
@@
Appendix B Contribution of nebular emission
We use the Yggdrasil models (Zackrisson et al., 2011) to estimate the possible contribution of nebular emission in rest-frame UV. As shown in Fig. 7, nebular emission is strongest for very young star-forming clouds, reaching up to $\sim 40\%$ of the stellar luminosity in moderately metal-enriched environments with $Z\geq 0.0004$. However, it decays with time and becomes negligible ($\lesssim 10\%$ of the stellar contribution) shortly ($\sim$ a few Myr $\ll t_{\rm SF}$) after the starburst. Besides, for continuous star formation with a constant SFR over the past $t_{\rm SF}=100~{}(10)\rm~{}Myr$, the nebular luminosity is below $10~{}(20)\%$ of the stellar luminosity. Note that Zackrisson et al. (2011) only consider non-rotatings stars without CHE and adopt a Kroupa (2001) IMF in the mass range $m_{\star}\in[0.1,100]~{}\rm M_{\odot}$ for $Z\geq 0.0004$. The nebular emission is stronger at lower metallicities and when the stellar ionizing radiation is enhanced. In the Yggdrasil models of metal-free clouds hosting non-rotating Pop III stars, nebular emission can dominate the rest-frame UV luminosity in the first few Myr and remain comparable to stellar emission for up to $\sim 20$ Myr after a starburst or continuous star formation with $t_{\rm SF}\sim 10-100$ Myr. Here, the nebular contribution is approximately proportional to the stellar ionizing flux. If this trend holds for moderately metal-enriched ($Z\gtrsim 0.0004$) galaxies most relevant for JWST observations, nebular emission can become comparable to stellar emission in rest-frame UV under CHE, since the production of ionizing photons is boosted by up to a factor of a few by CHE (Sibony et al., 2022, Liu et al. in prep.).
Figure 7: Ratio between nebular and stellar luminosities in rest-frame UV ($1500\pm 50~{}\rm\AA$) from an instantaneous starburst and continuous star formation (starting at $t=0$) with a constant rate over timescales $t_{\rm SF}=10$, 30, and 100 Myr (clockwise) for the Yggdrasil models (Zackrisson et al., 2011), assuming no leakage of ionizing photons (optimistic nebular emission). Among these are two models for Pop III stars with top-heavy IMFs: a Salpeter-like power-law ($dN/dm_{\star}\propto m_{\star}^{-2.35}$) in the range $m_{\star}\in[50,500]~{}\rm M_{\odot}$ (Pop III.1, solid) and a log-normal distribution with characteristic mass $m_{\star,\rm c}=10~{}\rm M_{\odot}$ and dispersion (in $\ln m_{\star}$) $\sigma=1$ for $m_{\star}\in[1,500]\ \rm M_{\odot}$ (Pop III.2, dashed), one Pop III model with a Kroupa (2001) IMF in $m_{\star}\in[0.1,100]\ \rm M_{\odot}$ (dash-dotted), three Pop II models with the same Kroupa (2001) IMF and different metallicities: $Z=0.0004$ (dotted), 0.004 (long-dashed), 0.008 (dash-dot-dotted), and a Population I (Pop I, densely-dashed) model with the same Kroupa (2001) IMF and $Z=0.02$. For the cases with continuous star formation, we are mainly interested in the results at $t=t_{\rm SF}$ (labeled with vertical lines), which denotes the moment of observation, i.e., when the galaxy has been forming stars for a duration of $t_{\rm SF}$.